# AnyPlace: Learning Generalized Object Placement for Robot Manipulation

### v1

@@ Abstract
Object placement in robotic tasks is inherently challenging due to the diversity of object geometries and placement configurations. To address this, we propose AnyPlace, a two-stage method trained entirely on synthetic data, capable of predicting a wide range of feasible placement poses for real-world tasks. Our key insight is that by leveraging a Vision-Language Model (VLM) to identify rough placement locations, we focus only on the relevant regions for local placement, which enables us to train the low-level placement-pose-prediction model to capture diverse placements efficiently. For training, we generate a fully synthetic dataset of randomly generated objects in different placement configurations (insertion, stacking, hanging) and train local placement-prediction models. We conduct extensive evaluations in simulation, demonstrating that our method outperforms baselines in terms of success rate, coverage of possible placement modes, and precision. In real-world experiments, we show how our approach directly transfers models trained purely on synthetic data to the real world, where it successfully performs placements in scenarios where other models struggle – such as with varying object geometries, diverse placement modes, and achieving high precision for fine placement. More at: any-place.github.io.
{strip}[Uncaptioned image]
@@
Figure 1: Execution of the AnyPlace approach by the robot. (1) Given a language description of a placement task, the robot first captures an RGBD image of the scene using its eye-in-hand camera. (2) A large segmentation model and a Vision Language Model (VLM) are then used to segment objects and suggest possible placement locations. (3) Multiple placement poses are predicted for objects around these suggested locations. (4) The robot is then able to insert a vial into different holes on the vial plate.
I Introduction
@@
Placing objects is a fundamental task that humans perform effortlessly in daily life, from setting items on a table to inserting cables into sockets. On the other hand, enabling a robot to perform such tasks can often be highly challenging. The challenges arise from the various constraints of different placement tasks and the difficulty of generalizing to unseen objects. Existing methods are often task-specific, using a large number of demonstrations for a single placement task, such as hanging objects on racks [28], with the hope that the robot can generalize to unseen objects. Alternatively, few-shot approaches focus on learning object placement with only a few demonstrations, aiming for the model to replicate the same placement operation across random initial configurations of similar objects and setups.
Learning diverse object placements presents significant challenges for existing models, primarily due to the difficulty of generalization. Generalization can be categorized into two aspects: object-level and task-level. Object-level generalization focuses on developing robust representations of various objects and placement configurations, enabling the model to handle unseen objects effectively. Task-level generalization involves the model’s ability to predict diverse placement configurations from given pointclouds. Additionally, predicting multimodal placement outputs—encompassing a range of valid locations and modes—remains challenging, particularly when multiple feasible solutions exist. For instance, a robot inserts a vial into one of many empty slots on a vial plate. Existing methods [26, 3, 27, 21, 24, 23, 9, 6, 15, 14] often operate within a few-shot learning framework, where they learn specific placements involving two particular objects, resulting in poor generalization to different tasks and object types. In contrast, RPDiff [28] has trained separate models for three placing scenarios (e.g. cup on a rack, book on a shelf, can stacking) using a relatively larger dataset, demonstrating some level of in-class object-level generalization, but its generalization to new tasks has not been studied.
@@
In this work, we address generalizable object placement that is robust to different objects and capable of predicting diverse placement poses across various tasks, similar to general-purpose object grasping methods. To achieve this, we have developed a fully synthetic dataset that contains 1,489 generated objects and captures three common placement configurations: inserting, stacking, and hanging. Based on this dataset, we develop a placement prediction pipeline that consists of a high-level placement position proposal module and a low-level placement pose prediction model. For the high-level module, we leverage a large segmentation model SAM-2 [22] to segment objects of interest and prompt the Molmo VLM [4] to propose all possible placement locations. Only the region around the proposed placement is fed into the low-level pose prediction model. By first providing a rough region for the low-level module to focus on, the model can effectively generalize across objects, learn their geometry, and capture diverse placement configurations. For the low-level pose prediction models, we build upon diffusion to predict precise and multimodal placement poses. We demonstrate the effectiveness of our methods across different placement tasks in simulation, where we significantly outperform baseline models in terms of pick-and-place success rate and placement location coverage. In real-world evaluation, our model achieves an 80% success rate on the vial insertion task, demonstrating robustness to noisy data and generalization to unseen objects.
To summarize, the key contributions of our work are:
@@
    We propose a novel object placement approach that leverages a VLM to reason about potential placement locations and a low-level pose prediction model to predict placement poses based solely on the region of interest.
@@
    We developed a fully synthetic dataset containing thousands of generated objects capturing a wide range of local placement configurations. This is crucial in enabling us to train models exclusively on synthetic data and be able to use them for a wide range of real-world placement tasks.
@@
    We show that reducing the problem to only local prediction allows us to improve performance with respect to baseline methods in terms of success rate, precision, and mode coverage. It also crucially makes it possible for a model trained on a fully synthetic dataset to show generalization that enables it to directly be deployed to the real world on a variety of placement tasks.
II Related Work
Figure 2: Overview of the AnyPlace placement pose prediction approach. (1) High-level Placement Location Prediction: Given an input language description and an RGBD image, we leverage a VLM and a segmentation model to extract objects of interest. Next, we prompt the VLM again to propose possible placement locations. Using camera parameters, we reproject the depth map into 3D and crop the region of interest centered on the proposed placement location. Full pointclouds of the objects to be placed, along with the cropped regions of the placement locations, are then fed into our low-level pose prediction model to output precise relative transformations for object placement. (2) Low-level Placement Pose Prediction: We use self-attention to extract pointcloud features and cross-attention to capture cross-object features. For the diffusion decoder, the diffusion timestep is encoded within the MLP layers before the model outputs the relative transformation.
II-A Object pick and place in robot manipulation
@@
The problem of robot pick-and-place is typically formulated in two ways: object rearrangement and direct end-effector pose prediction. In object rearrangement, the goal is to train a model to predict the relative transformation of the object from its initial pose to its final placement pose. Assuming the grasping pose is generated by existing models, the final placing pose of the end-effector is calculated by multiplying the predicted relative transformation with the grasping pose. In this setting, many of the works focus on predicting explicit task-relevant features of both objects and then solving for the relative pose through optimization or regression. Specifically, the Neural Descriptor Fields (NDF) series of papers [26, 3, 27] learn the occupancy field of pointclouds as a representation. Then, a set of predefined keypoints attached to the placement object is used to interact with and query the occupancy feature field of the target object. By matching the queried features at each point to features collected during demonstrations, the optimal object transformation is determined. TaxPose [21] leverages transformer-based cross-attention to predict corresponding points between two objects and use differentiable singular value decomposition (SVD) to solve for the relative transformation. To guarantee the placement pose prediction model is robust to SE(3) transformations, i.e., SE(3)-equivariant, methods [24, 23, 9] explicitly predict perpoint type-0 and type-1 features for object point clouds and then solve an optimization problem to align these features into specific configurations based on demonstrations. All of these methods operate in a few-shot setting and can predict a single placement pose given two objects. It is crucial for models to capture and predict a distribution of placement poses, as not every placement pose is realizable by a robot due to its kinematic constraints. RPDiff [28], by contrast, trains a transformer with a diffusion mechanism on a large dataset, gradually denoising the object placement pose. However, their experiments reveal that the coverage of possible placement locations is incomplete. The fixed-size cropping mechanism used during diffusion may also struggle to generalize to objects of varying sizes. Additionally, a recent study [5] samples multiple stable placements in a simulation and employs a VLM to select the appropriate mode based on a language query. While these modes are discrete, each mode allows for the rotation of objects along their axis of symmetry, resulting in valid placement poses that form a continuous distribution.
@@
An alternative approach to the pick-and-place task is predicting the robot’s end-effector pose directly. M2T2 [33], and Pick2Place [12] focus on planar object placement in cluttered scenes. M2T2 [33] employs a multi-task transformer with separate decoders to predict grasp poses and placement location affordance maps for each discrete bin of rotation. Other works, like Pick2Place [12], concentrate on predicting key end-effector poses to accomplish specific tasks. RVT [10] and RVT-2 [11] also utilizes a transformer, leveraging multiview RGB images of the scene to predict heatmaps for the robot’s next end-effector location. Coarse-to-fine Q-attention [16], on the other hand, leverages the scene’s voxel to identify the most interesting spatial point at the current resolution. This point becomes the voxel centroid for the next refinement step, enabling the model to gather more accurate 3D information.
II-B Multimodal prediction in robot manipulation
Consider a mug that can be hung on a rack in various poses at different pegs. Multimodality of placement poses is common in manipulation. Recently advancements in behavior cloning methods, such as the IBC energy-based model [8] and Diffusion Policy [2], have attempted to capture this multimodality. For diffusion policy specifically, it learns to denoise action trajectories starting from a random noise distribution, enabling the model to capture the underlying distribution of possible solutions. Another approach to modeling these distributions is through Variational Autoencoders (VAEs). BeT [25] and VQ-Bet [17], for instance, trained a VQ-VAE [30] to discretize and encode continuous actions into latent representations, making it easier to handle multimodal and high-dimensional behavior data. To train the entire policy based on image input, they employed a transformer-based model alongside the VAE decoder to predict sequences of actions. The authors demonstrated that their approach outperformed the Diffusion Policy [2] on common benchmarks.
III AnyPlace: Generalized Object Placement
@@
To enable a robot to execute diverse object placements in a scene, we propose breaking the placement pose prediction problem into two subtasks: a high-level placement location proposal task and a low-level fine-grained placement pose prediction task. For the high-level task, we incorporate a vision-language model, trained to output 2D keypoint locations in an image based on a given text prompt. A small local region around the candidate placement location can then be extracted and forwarded to the low-level pose-prediction model. This simplifies the low-level pose-prediction problem significantly by having a much smaller pointcloud as input and helps with generalization as any features outside of the local region do not influence the prediction. This allows us to focus on a limited set of general placement types and utilize a fully synthetic dataset, but have the final model be effective in a broad range of real-world placement tasks.
@@
Problem setup. We formulate the object placement task to predict relative transformations. Specifically, given an input tuple $\{D,I\}$, where $D$ represents the language description of the placement task and $I$ is an RGBD image of the scene, our goal is to predict a set of rigid transformations $\{T_{n}\}_{n=1}^{N}\subset\mathrm{SE}(3)$ that move the target object $C$ from its current position to all viable placement locations on the base object $B$ that satisfy language conditioning $D$. Assuming the grasping poses $T_{\text{grasp}}$ are provided by a grasp prediction model, the final end-effector pose $T_{\text{place}}$ can be computed using the predicted relative transformation between the initial object pose and its final placement pose as $T_{\text{place}}=T_{n}T_{\text{pick}}.$
III-A VLM-guided placement location prediction
@@
When objects have multiple potential placement locations, existing models often struggle to capture all these discrete positions when trained in an end-to-end manner. To address this, we propose leveraging recent advancements in VLMs, which have demonstrated strong capabilities in localizing points and regions within images based on language descriptions, to directly identify placement locations. Specifically, given a language description of the placement task $D$ and a RGBD image $I$, we aim to extract the pointcloud of the target object $\mathcal{P}_{\text{c}}$ and a local region of interest of the base object $\mathcal{P}_{\text{b\_crop}}$, where $\mathcal{P}_{\text{c}},\mathcal{P}_{\text{b\_crop}}\in\mathbb{R}^{N\times 3}$. We utilize Molmo [4] to detect all potential placement locations as keypoints in image space, such as specifying all positions where a vial can be inserted into a vial plate. This approach enables the low-level pose prediction model to focus on learning the different placement configurations of two objects and predicting the placement pose without needing to explicitly capture multiple placement locations.
Figure 3: Language Prompts and VLM Output Visualization in the Real-World Evaluation. In the real-world setting, we prompt the VLM to predict placement locations on a variety of different objects. Combined with our low-level pose prediction model, we demonstrate strong generalization and flexibility in real-world experiments. The red dot on each image is the predicted pixel location from the Molmo VLM.
First, we extract segmentation masks for the target object $C$ and the base object $B$. To achieve this, we first query the VLM to get point locations in the image for each object, then pass those points to the segmentation model to get segmentation masks. This allows us to have complex language conditioning in our object selection (e.g., ”blue vial”, ”vial rack in front of the scale”, etc.). Next, we query the VLM again on the base object $B$ to find all the discrete modes for the placement (Figure 3). For each identified point in the image, we extract the local region in the pointcloud and use that as input to our pose prediction model.
@@
Empirically, we found that explicitly identifying placement modes, rather than relying on models to explore placements across the entire object, is more reliable and practical when handling diverse objects with multiple possible placement poses. Additionally, since our high-level module is built on a general-purpose VLM, the system can handle diverse placements and perform complex language conditioning. In particular, this enables us to have complex language conditioning in: (1) selecting which object to place; (2) selecting which object to place it onto; and (3) selecting where on that object to perform the placement. It also enables our method to function effectively even when the object being placed onto is large or difficult to segment. This is because the module always outputs a small local region around the potential placement location, regardless of the object’s size.
III-B Fine-grained placement pose prediction

### v2

# AnyPlace: Learning Generalizable Object Placement for Robot Manipulation

@@ Abstract
Object placement in robotic tasks is inherently challenging due to the diversity of object geometries and placement configurations. We address this with AnyPlace, a two-stage method trained entirely on synthetic data, capable of predicting a wide range of feasible placement poses for real-world tasks. Our key insight is that by leveraging a Vision-Language Model (VLM) to identify approximate placement locations, we can focus only on the relevant regions for precise local placement, which enables us to train the low-level placement-pose-prediction model to capture multimodal placements efficiently. For training, we generate a fully synthetic dataset comprising 13 categories of randomly generated objects in 5370 different placement poses across three configurations (insertion, stacking, hanging) and train local placement-prediction models. We extensively evaluate our method in high-fidelity simulation and show that it consistently outperforms baseline approaches across all three tasks in terms of success rate, coverage of placement modes, and precision. In real-world experiments, our method achieves an average success and coverage rate of 76% across three tasks, where most baseline methods fail completely. We further validate the generalization of our approach on 16 real-world placement tasks, demonstrating that models trained purely on synthetic data can be directly transferred to the real world in a zero-shot setting. More at: any-place.github.io.
@@
Figure 1: Execution of the AnyPlace approach by the robot. (1) Given a language description of a placement task, the robot first captures an RGBD image of the scene using its eye-in-hand camera. (2) A segmentation model and a VLM are used to segment objects and suggest possible placement locations. (3) Multiple placement poses are predicted for objects around suggested locations. (4) The robot realizes placement into any of the predicted poses. (A-O) AnyPlace shows generalization and robustness in predicting placement poses across 16 tasks in the real world, despite being trained purely on a small synthetic dataset.
Keywords: Pick and Place, Robot Manipulation, Synthetic Dataset
1 Introduction
@@
Placing objects is a fundamental task that humans perform effortlessly in daily life, from setting items on a table to inserting cables into sockets. On the other hand, enabling a robot to perform such tasks can often be highly challenging. The challenges arise from the various constraints of different placement tasks and the difficulty of generalizing to unseen objects. Additionally, predicting multimodal placement outputs, which encompass a range of valid locations and modes, remains challenging, particularly when multiple feasible solutions exist. Existing methods are often task-specific, using a large number of demonstrations for a single placement task, such as hanging objects on racks [1], hoping that the robot can generalize to unseen objects. Alternatively, few-shot approaches [2, 3, 4, 5, 6, 7, 8, 9, 10, 11] focus on learning object placement with a few demonstrations, aiming for the model to replicate the same placement operation across random initial configurations of similar objects and setups. However, both struggle with generalization and scalability.
@@
We reframe the placement problem as a pairwise shape mating [12], allowing us to unify multiple placement tasks into a single learning objective. In this work, we address generalizable object placement that is robust to different objects and capable of predicting diverse and precise placement poses across various tasks, such as placing in open areas, inserting in fine slots, and hanging. We have developed a fully synthetic dataset that captures three common placement configurations: inserting, stacking, and hanging. Furthermore, we develop a placement prediction algorithm that consists of a high-level placement position proposal module and a low-level placement pose prediction model. We use a spatial VLM to propose all possible placement locations and extract the local point cloud regions based on them. By providing a coarse region for the low-level module to focus on, the model can effectively generalize across objects, learn their geometry, and capture diverse placement configurations. We use a diffusion model for fine pose prediction, conditioned on a small local point cloud region, which enables precise and multimodal placement pose predictions. We demonstrate the effectiveness of our methods across a range of placement tasks in simulation, as well as diverse real-world tasks, outperforming the baseline models in terms of success rate and coverage in both cases. The key contributions of our work are:
@@
    We propose a novel object placement approach that leverages a VLM to reason about potential placement locations and a low-level pose prediction model to predict placement poses based solely on the region of interest. We show that this coarse-to-fine mechanism allows us to significantly improve performance with respect to baseline methods in terms of success rate, precision, and placement mode coverage.
@@
    We develop a data generation pipeline and build a fully synthetic dataset containing thousands of generated objects and capturing a wide range of local placement configurations.
@@
    We demonstrate our approach generalizes to the real world on 16 placement tasks with novel objects and varying degrees of precision requirements.
2 Related Work
The problem of robot pick-and-place is typically formulated in two ways: object rearrangement and direct end-effector pose prediction.
@@ 2 Related Work
Object Rearrangement In object rearrangement, the goal is to train a model to predict the relative transformation of the object from its initial pose to its final placement pose. In this setting, many of the works focus on predicting explicit task-relevant features of both objects and then solving for the relative pose through optimization or regression. Specifically, the Neural Descriptor Fields (NDF) series of papers [2, 3, 4] learn the occupancy field of point clouds as a representation. A fixed set of keypoints on the placement object queries features from the target’s field, and the best transformation is estimated by matching these to demonstration features. TaxPose [5] leverages transformer-based cross-attention to predict corresponding points between two objects and uses differentiable singular value decomposition (SVD) to solve for the relative transformation. To guarantee the placement pose prediction model is robust to SE(3) transformations, i.e., SE(3)-equivariant, methods [6, 7, 8] explicitly predict per point type-0 and type-1 features for object point clouds and then solve an optimization problem to align these features into specific configurations based on demonstrations. All of these methods operate in a few-shot setting and can predict a single placement pose given two objects. It is crucial for models to capture and predict a distribution of placement poses, as not every placement pose is realizable by a robot due to its kinematic constraints. RPDiff [1], by contrast, trains a transformer with a diffusion mechanism on a large dataset, gradually denoising the object placement pose. However, their experiments reveal that the coverage of possible placement locations is incomplete. The fixed-size cropping mechanism used during diffusion may also struggle to generalize to objects of varying sizes. Additionally, a recent study [13] samples multiple stable placements in a simulation and employs a VLM to select the appropriate mode based on a language query. While these modes are discrete, each mode allows for the rotation of objects along their axis of symmetry, resulting in valid placement poses that form a continuous distribution.
@@
Direct Pick and Place End-effector Pose Prediction An alternative approach to the pick-and-place task is predicting the robot’s end-effector pose directly. M2T2 [14] and Pick2Place [15] focus on planar object placement in cluttered scenes. M2T2 [14] employs a multi-task transformer with separate decoders to predict grasp poses and placement location affordance maps for each discrete bin of rotation. Other works, like Pick2Place [15], concentrate on predicting key end-effector poses to accomplish specific tasks. RVT [16] and RVT-2 [17] also utilize a transformer, leveraging multiview RGB images of the scene to predict heatmaps for the robot’s next end-effector location. Coarse-to-fine Q-attention [18], on the other hand, leverages the scene’s voxel to identify the most interesting spatial point at the current resolution. This point becomes the voxel centroid for the next refinement step, enabling the model to gather more accurate 3D information. Another line of work on contemporary VLMs and vision foundation models (VFMs) [19] finds that they often lack reliable spatial reasoning abilities, limiting their effectiveness in fine-grained manipulation tasks such as peg-in-hole insertion, where precise object placement is critical.
3 AnyPlace: Generalizable Object Placement
@@
To enable a robot to execute diverse object placements in a scene, we propose decomposing the placement pose prediction problem into two stages: a high-level coarse placement location proposal stage and a low-level fine placement pose prediction stage. For the high-level task, we incorporate a vision-language model, trained to output 2D keypoint locations in an image based on a given text prompt. A small local region around the candidate placement location can then be extracted for the low-level pose-prediction model. This simplifies the low-level pose prediction problem significantly by using a much smaller point cloud as input and improves generalization overall, as features outside the local region do not influence the prediction. This allows us to focus on a limited set of general placement types and utilize a fully synthetic dataset, but have the final model be effective in a broad range of real-world placement tasks. Additionally, the high-level prediction stage enables the identification of multiple placement modes.
@@
Problem setup. We formulate the object placement task as predicting relative transformations. Specifically, given an input tuple $\{D,I\}$, where $D$ represents the language description of the placement task and $I$ is an RGBD image of the scene, our goal is to predict a set of rigid transformations $\{T_{n}\}_{n=1}^{N}\subset\mathrm{SE}(3)$ that move the target object $C$ from its current position to all viable placement locations on the base object $B$ that satisfy language conditioning $D$. Assuming the grasping poses $T_{\text{grasp}}$ are provided by a grasp prediction model, the final end-effector pose $T_{\text{place}}$ can be computed using the predicted relative transformation between the initial object pose and its final placement pose as $T_{\text{place}}=T_{n}T_{\text{pick}}.$
Figure 2: Overview of the AnyPlace placement pose prediction approach. (1) High-level Coarse Placement Location Prediction: Given an input language description and an RGBD image, we leverage a VLM and a segmentation model to extract objects of interest. We then prompt the VLM to propose possible placement locations and crop the region of interest centered on the proposed points. The resulting point clouds are fed into the low-level model. (2) Low-level Fine Placement Pose Prediction: We use a transformer architecture, with a diffusion decoder to predict relative transformation of the point clouds to realize placement.
3.1 VLM-guided coarse placement location prediction
@@
When multiple potential placement locations exist within a task, existing models often struggle to capture all possibilities. To address this, we propose leveraging spatial VLMs, which have demonstrated strong capabilities in localizing points and regions within images based on language descriptions, to directly identify placement locations. Specifically, given a language description of the placement task $D$ and a RGBD image $I$, we extract the point cloud of the target object $\mathcal{P}_{\text{c}}$, and for each identified placement point in the image, we extract the local region of the base object $\mathcal{P}_{\text{b\_crop}}$ guided by the 3D bounding box of the target object, where $\mathcal{P}_{\text{c}},\mathcal{P}_{\text{b\_crop}}\in\mathbb{R}^{N\times 3}$. These point clouds are then used as input to the pose prediction model.
@@
This approach enables the low-level pose prediction model to focus on learning different placement configurations of two objects and predicting placement poses. Also, explicitly identifying placement modes, rather than relying on models to explore placements across the entire object, is more reliable and practical when handling diverse objects with multiple possible placement poses. Since our high-level module is built on a general-purpose VLM, the system can also handle diverse placements and perform complex language conditioning. We utilize Molmo [20] to detect all potential placement locations as keypoints in image space, such as specifying all positions where a vial can be inserted into a vial plate. However, our approach is not specific to one VLM, and other models with capabilities to give point locations in an image could be used [21]. (See Appendix A for language prompts.)
3.2 Fine-grained placement pose prediction